# Planck stars, White Holes, Remnants and Planck-mass quasi-particles The quantum gravity phase in black holes’ evolution and its manifestations

### v3

@@ II.6 White holes
Before addressing the dissipative aspects of the life of a hole, let us pause to clarify a few properties of the white holes. White holes, like black holes, are exact solutions of the Einstein’s equation. Precisely like black holes, they have long been expected not to play any role in our universe, by a majority of physicists. The situation has changed for black holes in the last decades. It has not yet equally changed for white holes, but it might soon.
@@
This is clear from the usual Schwarzschild line element, which is symmetric under time reversal, and therefore describes equally well the exterior of a black hole and the exterior of a white hole. Equivalently, zone II of the maximal extension of the Schwarzschild solution is equally the outside of a black hole and the outside of a white hole (see Fig. 1, Left). Analogous considerations hold for the Kerr solution. The continuation of the external metric of a stationary Kerr or Schwarzschild spacetime inside the radius $r=2m+\epsilon$ contains both a trapped region (a black hole) ad an anti-trapped region (a white hole).
Figure 13: Left: in the extended Schwarzschild spacetime, which is stationary, the (light grey) region outside $r=2m+\epsilon$ (dotted line) is equally the outside of a black and a white hole. Center: A collapsing matter distribution (dark grey) replaces the white hole region (“WH”) in the non-stationary collapse metric. Right: The time revered process. The difference between the last two can only be detected looking a the past, or the future.
What distinguishes then the objects we call ‘black holes’ from ‘white holes’? The objects in the sky we call ‘black holes’ are described by a stationary metric only approximately and for a limited time. We expect that (at least) in the past their metric was definitely non-stationary, and they were produced by gravitational collapse. The energy contained inside $r=2m+\epsilon$ was less than $m$ in the past and the continuation of the metric inside this radius contains the trapped region, but not the anti-trapped region, which is instead replaced by the region describing the collapsing star. Therefore seen from the outside a ‘black hole’ (as opposite to a ‘white hole’) is only characterised by the fact that in the past it does not have an anti-trapped region (see Fig. 1, Center). Viceversa, a white hole is an object that from the exterior and for a finite time is indistinguishable from a black hole, but in the future ceases to be stationary, the amount of energy inside $r=2m+\epsilon$ decreases, and there is no trapped region in the future (see Fig.1, Right). All this regards classical gravity only.
@@ III.1 Black hole lifetime
The lifetime $\tau_{BH}$ of a black hole is known by Hawking radiation theory. It can be estimated as follows. The Hawking radiation is thermal, namely it has Planck spectrum. The peak of the spectrum is on a wavelength $\lambda$ that is determined by the only length parameter in the problem: the Schwarzschild radius $2Gm/c^{2}$. Therefore the radiation is mostly composed by quanta of frequency
@@
The evaporation shrinks the horizon, bringing it close to Planckian size, where the black-to-white transition has high probability to happen. The white hole generated by the process has then an horizon of Planckian size. (This lifetime can be shorter if quantum gravity fluctuations trigger an earlier tunnelling. This could be as early as $\tau_{BH}\sim m_{o}^{2}$ as suggested in Haggard2014 ; Haggard20162 . Here we only focus on the possibility that the transition happens at the end of the evaporation.)
@@ III.2 How big is the interior of a black hole?
In Minkowski spacetime, we say that a space-like 2-sphere $S$ of radius $r$ encloses a space-like ball of volume $\frac{4}{3}\pi r^{3}$. As a 3d surface in Minkowski space, this ball is characterized by two properties: it is linear, and it is the surfaces that maximizes the volume among those bounded by $S$. Any deformation, in fact, is time-like and decreases the volume.
@@ III.3 Remnants and their lifetime
The argument of the previous section suggest that at the end of the evaporation a black hole undergoes a quantum transition to a white hole with a Planckian-size horizon and a vast interior.
The possibility of remnants with a similar structure was considered in the 1990’s Giddings1992b . What was not realized at that time is that classical GR does in fact predict the existence of objects with the these properties: white holes; and quantum theory could account for their formation at the end of the evaporation and (as we shall see below) for their stability.
Before studying more in detail the quantum structure of these remnants, let us ask what could be their lifetime.
@@
The second argument concerns the information trapped inside the hole. The Hawking radiation is genuinely thermal: it is describe by a quantum state which is not pure: it is a density matrix. This is because the Hawking radiation that escapes to infinity forms together with negative radiation that falls inside the hole. The two are entangled, therefore the part escaping to infinity has von Newman entropy. The total entropy of the Hawking radiation is of the order of the Bekenstein Hawking entropy
@@
where $A\sim m_{o}^{2}$ is the initial area of the black hole, before the beginning of the evaporation. Some scientists expect that the total entropy of the Hawking radiation start decreasing at Page time because late Hawking quanta are correlated with early ones. We are not convinced by this expectation for reasons that we explain in detail later on. Hence an amount of information of the order $S_{BH}$ is trapped inside the hole during the quantum transition. This must be emitted later in the form or radiation by the remnant. The remnant has total energy of the order of the Planck mass, namely of order unit in natural units. It must therefore be able to emit a large amount of information by emitting very small energy.
@@
Notice that the state of the black hole at some given (advanced) time, which determines its full future evolution, is not specified by its sole current mass $m$, but also by the internal geometry, which in turn is determined by the initial mass of the black hole.
@@
These are time scales for a distant observer. If $m_{o}$ is large, the process is very long in the time of a distant observer. But it is extremely short (of order $m_{o}$, which is the time light takes to cross the radius of the star) if measured on the bouncing matter distribution itself. The huge difference is due to the extreme gravitational time dilation Haggard2014 . Time slows down near high density mass. An observer (capable of resisting the tidal forces) landing on a Planck matter distribution will find herself nearly immediately in the distant future, at the time where the black hole ends its evaporation. The proper lifetime of a Planck matter distribution is short: from its own perspective, the matter distribution is essentially a bounce. A black hole is a shortcut to the distant future.
This is a compelling scenario, but it still need to be corrected.
@@ III.4 Instability
To see this, consider a black hole of mass $m$, formed by a collapsed star. In the unperturbed solution, no energy reaches future null infinity (here we are considering classical physics: no Hawking radiation). Now, consider a perturbation of this spacetime in which there is a small spherically symmetric pulse of null radiation with total energy $\epsilon$ arriving on null infinity. Here $\epsilon$ can be arbitrarily small. In order to arrive to null infinity, this radiation must be emitted by the surface of the matter distribution before this enters the horizon. Say it was emitted when the radius of the matter distribution was $2m+\delta$. But this must also be outside the Schwarzschild radius of the matter distribution plus the energy to be emitted, otherwise it could not reach null infinity either. Hence necessarily $2m+\delta>2m+\epsilon$ that is, we must have $\delta>\epsilon$. Now let’s $u_{\delta}$ the retarded time of the emission point. Necessarily, the energy pulse of energy $\epsilon$ must reach null infinity before $u_{\delta}$. Therefore, no perturbation with energy $\epsilon$ can reach null infinity after $u_{\delta}$. That is, we cannot perturb the the final data arbitrarily. However small is $\epsilon$ there are locations where it cannot get to.
@@
The same is true for a white hole, reversing the time direction. In the spacetime depicted in the Right panel, with some radiation, there is necessarily a dark spot in the incoming radiation from past null infinity. If we perturb this configuration, and add some incoming radiation in this dark spot, the evolution generically gives the spacetime of the Left panel.
Physically, what happens is that this radiation moves along the horizon, is blue shifted, can meet radiation coming out of the white hole and this is more mass that $m$ at a radius $r\sim 2m$: it is mass inside its Schwarzschild radius. At this point the region is trapped, and a black hole forms. Consequently the evolution of the perturbed initial conditions yields the spacetime of the Left, not the one of the Right: the white hole is unstable and decays into a black hole. This is the standard “instability of white holes” in classical GR.
@@ III.5 Planckian Remnants
But let nevertheless assume that a Planck-scale white hole is unstable. Its decay mode, as explained above, is into a Planck-scale black hole. Using the notation introduced above, this is the process
@@
In quantum mechanics, two such processes imply that a system free to loose energy into radiation will settle on its lowest available energy eigen-state, which will be a superposition of the two states, of the form
@@
Since the black and white hole are indistinguishable from the exterior, this superposition has no effect on the exterior. It is only a quantum spread of the geometry on and just around the horizon. In the next section we describe it more precisely.
Minimal energy, for the horizon, means minimal area. However, vanishing area means flat space, namely total dissipation. The transition to this state is likely to be strongly suppressed by the arguments in the last section. Intuitively, this is because it takes time to re-absorb the huge interior, and its entropic content. Hence, the system will settle for a long time on the lowest nonvanishing eigenstate of the area.
@@
That is, LQG predicts that black holes end up as long living particles with a mass of a few micro-grams.
Once a remnant has attended this minimal size, it is semi-stable. It can still radiate away its energy, and dissipate into flat space. But this last transition
@@
The existence of there objects contradicts the intuition of many particle-physics or string string theorists. This is because in flat space physics, a small volume with a small energy can contain only a finite number of different states. This is what prevents the blackbody UV catastrophe. Intuitively, to have many states we need many particles, and to have many particles with small total energy we need to have them with long wavelength, but this is not available in the volume contained within a small sphere, because this has a small volume. This is the basis of the intuition of many particle-physics and string theorists.
In general relativistic physics, however, these constraints do not hold, because on a curved geometry an arbitrarily large volume can be enclosed into an arbitrary small sphere. Hence an arbitrarily small horizon can enclose any number of different states, limited only by the maximal size $m_{o}$ of the parent black hole and by the finiteness of the time during which the black hole interior have had the opportunity to grow.
@@ III.6 Remnants and quantum spread of the extrinsic curvature
Classical black holes are described by semiclassical quantum states that are peaked on large values of $A$ and small values of $k$, with $\Delta A$ and $\Delta k$ satisfying the Heisenberg relations but all relative spreads being small. Now consider an isolated black hole that radiates away all its energy, but is stuck on the minimal non vanishing area because of the suppression of the transition to Minkowski due to its large interior. Then the system settles on an eigenstate of $A$. This is the common situation of all quantum system, left alone and free to equilibrate by radiating away energy: they settle on their minimal reachable energy eigenstate. An eigenstate of the area will have $k$ maximally spread. Therefore we expect Planck scale remnants to be in states where $k$ is maximally spread. Now consider a white hole. Its metric is easilly obtained time reversing the black hole metric, that is
@@ III.7 There is no information paradox
We close this part of the paper by discussing the vexata questio of the black hole information paradox, at the light of the above results. This Section follows closely Rovelli2017a and Rovelli2019a .
@@
The possibility that $N\gg N_{BH}$ is supported by the fact that according to classical GR the interaction between a black hole and its surroundings is entirely determined by what happen in the vicinity of the horizon, not by the interior. The number $N_{BH}$ is likely to count only horizon states that (before the transition) can be distinguishable from the exterior. These are “surface” states. On the other hand, $N$ counts also states that can be distinguished by local observables inside the horizon. These do not contribute to the thermodynamical entropy $S={A/4}$, but can contribute to the von Neumann entropy, which can therefore remain high even when $A$ shrinks.
These generic considerations can be sharpened. The fact that a blackhole can have more states than $N_{BH}$, in facts, follows from elementary considerations of causality.
@@
Figure 16: The (lowest part) of the conformal diagram of a gravitational collapse. The clear grey region is the object, the dotted line is the horizon, the thick upper line is the singularity, the dark upper region is where quantum gravity effect may become relevant (this region play no role in this paper.) The two Cauchy surfaces used in the paper are the dashed lines.
@@
Care is required in specifying what is meant here by ‘horizon’, since there are several such notions (event horizon, trapping horizon, apparent horizon, dynamical horizon…) which in this context may give tiny differences in location. For precision, by ‘horizon’ we mean here the event horizon, if this exist. If it doesn’t, we mean the boundary of the past of a late-time spacelike region lying outside the black hole (say outside the trapping region). With this definition, the horizon is light-like.
@@
Now consider the evolution of the quantum fields from $\Sigma_{1}$ to $\Sigma_{2}$. We are in a region far away from the singularity and therefore (assuming the black hole is large) from high curvature. Therefore we expects conventional quantum field theory to hold here, without strange quantum gravity effects, at least up to high energy scales. Since the horizon is light-like, $\Sigma_{1}^{in}$ is in the causal past of $\Sigma_{2}^{in}$. This implies that any local observable on $\Sigma_{1}^{in}$ is fully determined by observables on $\Sigma_{2}^{in}$. That is, if ${\cal A}_{i}$ is the local algebra of observables on $\Sigma_{i}^{in}$ then ${\cal A}_{1}$ is a subalgebra of ${\cal A}_{2}$:
@@
Therefore any state on ${\cal A}_{2}$ is also a state on ${\cal A}_{1}$ and if two such states can be distinguished by observables in ${\cal A}_{1}$ they certainly can be distinguished by observables in ${\cal A}_{2}$ as the first are included in the latest. Therefore the states that can be distinguished by ${\cal A}_{1}$ —which is to say: on $\Sigma_{1}^{in}$— can also be distinguished by ${\cal A}_{2}$ —which is to say: on $\Sigma_{2}^{in}$. Therefore the distinguishable states on $\Sigma_{1}^{in}$ are a subset of those in $\Sigma_{2}^{in}$. How many are them? Either there is an infinite number of them, or a finite number due to some high-energy (say Planckian) cut-off. If there is an infinite number of them, then immediately the number of states distinguishable from inside the black hole is larger that $N_{NB}$, which is finite. If there is a finite number of them, then the number $N_{2}$ of distinguishable states on $\Sigma_{2}^{in}$ must be equal or larger than the number $N_{2}$ of states distinguishable on $\Sigma_{1}^{in}$, because the second is a subset of the first. That is
@@
The physical interpretation of the conclusion is simple: the thermal behaviour of the black hole described by the Bekenstein-Hawking entropy $S=A/4$ is determined by the physics of the vicinity of the horizon.
A vivid manifestation of the fact that in classical GR the effect of a black hole on its surroundings is independent from the black hole interior is in the numerical simulations of black hole merging and radiation emission by oscillating black holes: in writing the numerical code, it is routine to cut away a region inside the (trapping) horizon: it is irrelevant for whatever happens outside! This is true in classical GR, and there is no compelling reason to suppose it to fail if quantum fields are around.
@@
In conclusion, if there are more states available in a black hole than $e^{A/4}$, then Page argument for the information loss paradox fails. States purifying the Hawking radiation are inside the hole even when the horizon shrinks. The information can later escape from the remnant emission (see below).

### v4

@@ II.6 White holes
Before addressing the dissipative aspects of the life of a hole, let us pause to clarify a few properties of the white holes. White holes, like black holes, are exact solutions of the Einstein’s equation. Precisely like black holes, they have long been expected to not play any role in our universe, by a majority of physicists. The situation has changed for black holes in the last decades. It has not yet equally changed for white holes, but it might soon.
@@
This is clear from the usual Schwarzschild line element, which is symmetric under time reversal, and therefore describes equally well the exterior of a black hole and the exterior of a white hole. Equivalently, zone II of the maximal extension of the Schwarzschild solution is equally the outside of a black hole and the outside of a white hole (see Fig. 1, Left). Analogous considerations hold for the Kerr solution. The continuation of the external metric of a stationary Kerr or Schwarzschild spacetime inside the radius $r=2m+\epsilon$ contains both a trapped region (a black hole) and an anti-trapped region (a white hole).
Figure 13: Left: in the extended Schwarzschild spacetime, which is stationary, the (light grey) region outside $r=2m+\epsilon$ (dotted line) is equally the outside of a black and a white hole. Center: A collapsing matter distribution (dark grey) replaces the white hole region (“WH”) in the non-stationary collapse metric. Right: The time-revered process. The difference between the last two can only be detected looking at the past, or the future.
What distinguishes then the objects we call ‘black holes’ from ‘white holes’? The objects in the sky we call ‘black holes’ are described by a stationary metric only approximately and for a limited time. We expect that (at least) in the past their metric was definitely non-stationary, and they were produced by gravitational collapse. The energy contained inside $r=2m+\epsilon$ was less than $m$ in the past and the continuation of the metric inside this radius contains the trapped region, but not the anti-trapped region, which is instead replaced by the region describing the collapsing star. Therefore seen from the outside a ‘black hole’ (as opposite to a ‘white hole’) is only characterized by the fact that in the past it does not have an anti-trapped region (see Fig. 1, Center). Vice versa, a white hole is an object that from the exterior and for a finite time is indistinguishable from a black hole, but in the future ceases to be stationary, the amount of energy inside $r=2m+\epsilon$ decreases, and there is no trapped region in the future (see Fig.1, Right). All this regards classical gravity only.
@@ III.1 Black hole lifetime
The lifetime $\tau_{BH}$ of a black hole is known by Hawking radiation theory. It can be estimated as follows. The Hawking radiation is thermal, namely it has a Planck spectrum. The peak of the spectrum is on a wavelength $\lambda$ that is determined by the only length parameter in the problem: the Schwarzschild radius $2Gm/c^{2}$. Therefore the radiation is mostly composed by quanta of frequency
@@
The evaporation shrinks the horizon, bringing it close to Planckian size, where the black-to-white transition has high probability to happen. The white hole generated by the process then has a horizon of Planckian size. (This lifetime can be shorter if quantum gravity fluctuations trigger an earlier tunnelling. This could be as early as $\tau_{BH}\sim m_{o}^{2}$ as suggested in Haggard2014 ; Haggard20162 . Here we only focus on the possibility that the transition happens at the end of the evaporation.)
@@ III.2 How big is the interior of a black hole?
In Minkowski spacetime, we say that a space-like 2-sphere $S$ of radius $r$ encloses a space-like ball of volume $\frac{4}{3}\pi r^{3}$. As a 3d surface in Minkowski space, this ball is characterized by two properties: it is linear, and it is the surface that maximizes the volume among those bounded by $S$. Any deformation, in fact, is time-like and decreases the volume.
@@ III.3 Remnants and their lifetime
The argument of the previous section suggests that at the end of the evaporation a black hole undergoes a quantum transition to a white hole with a Planckian-size horizon and a vast interior.
The possibility of remnants with a similar structure was considered in the 1990’s Giddings1992b . What was not realized at that time is that classical GR does in fact predict the existence of objects with these properties: white holes; and quantum theory could account for their formation at the end of the evaporation and (as we shall see below) for their stability.
Before studying in more detail the quantum structure of these remnants, let us ask what could be their lifetime.
@@
The second argument concerns the information trapped inside the hole. The Hawking radiation is genuinely thermal: it is describe by a quantum state which is not pure: it is a density matrix. This is because the Hawking radiation that escapes to infinity forms together with negative radiation that falls inside the hole. The two are entangled, therefore the part escaping to infinity has von Newman entropy. The total entropy of the Hawking radiation is of the order of the Bekenstein-Hawking entropy
@@
where $A\sim m_{o}^{2}$ is the initial area of the black hole, before the beginning of the evaporation. Some scientists expect that the total entropy of the Hawking radiation starts decreasing at Page time because late Hawking quanta are correlated with early ones. We are not convinced by this expectation for reasons that we explain in detail later on. Hence an amount of information of the order $S_{BH}$ is trapped inside the hole during the quantum transition. This must be emitted later in the form or radiation by the remnant. The remnant has total energy of the order of the Planck mass, namely of order unity in natural units. It must therefore be able to emit a large amount of information by emitting very small energy.
@@
Notice that the state of the black hole at some given (advanced) time, which determines its full future evolution, is not specified only by its current mass $m$, but also by the internal geometry, which in turn is determined by the initial mass of the black hole.
@@
These are time scales for a distant observer. If $m_{o}$ is large, the process is very long in the time of a distant observer. But it is extremely short (of order $m_{o}$, which is the time light takes to cross the radius of the star) if measured on the bouncing matter distribution itself. The huge difference is due to the extreme gravitational time dilation Haggard2014 . Time slows down near high density mass. An observer (if capable of resisting the tidal forces) landing on a Planck matter distribution will find herself nearly immediately in the distant future, at the time where the black hole ends its evaporation. The proper lifetime of a Planck matter distribution is short: from its own perspective, the matter distribution is essentially a bounce. A black hole is a shortcut to the distant future.
This is a compelling scenario, but it still needs to be corrected.
@@ III.4 Instability
To see this, consider a black hole of mass $m$, formed by a collapsed star. In the unperturbed solution, no energy reaches future null infinity (here we are considering classical physics: no Hawking radiation). Now, consider a perturbation of this spacetime in which there is a small spherically symmetric pulse of null radiation with total energy $\epsilon$ arriving at null infinity. Here $\epsilon$ can be arbitrarily small. In order to arrive at null infinity, this radiation must be emitted by the surface of the matter distribution before this enters the horizon. Say it was emitted when the radius of the matter distribution was $2m+\delta$. But this must also be outside the Schwarzschild radius of the matter distribution plus the energy to be emitted, otherwise it could not reach null infinity either. Hence necessarily $2m+\delta>2m+\epsilon$ that is, we must have $\delta>\epsilon$. Now let $u_{\delta}$ be the retarded time of the emission point. Necessarily, the energy pulse of energy $\epsilon$ must reach null infinity before $u_{\delta}$. Therefore, no perturbation with energy $\epsilon$ can reach null infinity after $u_{\delta}$. That is, we cannot perturb the final data arbitrarily. However small is $\epsilon$ there are locations where it cannot get to.
@@
The same is true for a white hole, reversing the time direction. In the spacetime depicted in the Right panel, with some radiation, there is necessarily a dark spot in the incoming radiation from past null infinity. If we perturb this configuration, and add some incoming radiation to this dark spot, the evolution generically gives the spacetime of the Left panel.
Physically, what happens is that this radiation moves along the horizon, is blue shifted, can meet radiation coming out of the white hole and this is more mass than $m$ at a radius $r\sim 2m$: it is mass inside its Schwarzschild radius. At this point the region is trapped, and a black hole forms. Consequently the evolution of the perturbed initial conditions yields the spacetime of the Left, not the one of the Right: the white hole is unstable and decays into a black hole. This is the standard “instability of white holes” in classical GR.
@@ III.5 Planckian Remnants
But let’s nevertheless assume that a Planck-scale white hole is unstable. Its decay mode, as explained above, is into a Planck-scale black hole. Using the notation introduced above, this is the process
@@
In quantum mechanics, two such processes imply that a system free to lose energy into radiation will settle on its lowest available energy eigen-state, which will be a superposition of the two states, of the form
@@
Since the black and white holes are indistinguishable from the exterior, this superposition has no effect on the exterior. It is only a quantum spread of the geometry on and just around the horizon. In the next section we describe it more precisely.
Minimal energy, for the horizon, means minimal area. However, vanishing area means flat space, namely total dissipation. The transition to this state is likely to be strongly suppressed by the arguments in the last section. Intuitively, this is because it takes time to re-absorb the huge interior, and its entropic content. Hence, the system will settle and remain on the lowest nonvanishing eigenstate of the area for a long time.
@@
That is, LQG predicts that black holes end up as long-living particles with a mass of a few micro-grams.
Once a remnant has attained this minimal size, it is semi-stable. It can still radiate away its energy, and dissipate into flat space. But this last transition
@@
The existence of these objects contradicts the intuition of many particle-physics or string string theorists. This is because in flat space physics, a small volume with a small energy can contain only a finite number of different states. This is what prevents the blackbody UV catastrophe. Intuitively, to have many states we need many particles, and to have many particles with small total energy we need to have them with long wavelength, but this is not available in the volume contained within a small sphere, because this has a small volume. This is the basis of the intuition of many particle-physics and string theorists.
In general-relativistic physics, however, these constraints do not hold, because on a curved geometry an arbitrarily large volume can be enclosed into an arbitrarily small sphere. Hence an arbitrarily small horizon can enclose any number of different states, limited only by the maximal size $m_{o}$ of the parent black hole and by the finiteness of the time during which the black hole interior has had the opportunity to grow.
@@ III.6 Remnants and quantum spread of the extrinsic curvature
Classical black holes are described by semiclassical quantum states that are peaked on large values of $A$ and small values of $k$, with $\Delta A$ and $\Delta k$ satisfying the Heisenberg relation but all relative spreads being small. Now consider an isolated black hole that radiates away all its energy, but is stuck on the minimal non-vanishing area because of the suppression of the transition to Minkowski due to its large interior. Then the system settles on an eigenstate of $A$. This is the common situation of all quantum systems, left alone and free to equilibrate by radiating away energy: they settle on their minimal reachable energy eigenstate. An eigenstate of the area will have $k$ maximally spread. Therefore we expect Planck scale remnants to be in states where $k$ is maximally spread. Now consider a white hole. Its metric is easily obtained time reversing the black hole metric, that is
@@ III.7 There is no information paradox
We close this part of the paper by discussing the vexata questio of the black hole information paradox, in light of the above results. This Section follows closely Rovelli2017a and Rovelli2019a .
@@
The possibility that $N\gg N_{BH}$ is supported by the fact that according to classical GR the interaction between a black hole and its surroundings is entirely determined by what happens in the vicinity of the horizon, not by the interior. The number $N_{BH}$ is likely to count only horizon states that (before the transition) can be distinguishable from the exterior. These are “surface” states. On the other hand, $N$ counts also states that can be distinguished by local observables inside the horizon. These do not contribute to the thermodynamical entropy $S={A/4}$, but can contribute to the von Neumann entropy, which can therefore remain high even when $A$ shrinks.
These generic considerations can be sharpened. The fact that a blackhole can have more states than $N_{BH}$, in fact, follows from elementary considerations of causality.
@@
Figure 16: The (lowest part) of the conformal diagram of a gravitational collapse. The clear grey region is the object, the dotted line is the horizon, the thick upper line is the singularity, the dark upper region is where quantum gravity effects may become relevant (this region plays no role in this paper.) The two Cauchy surfaces used in the paper are the dashed lines.
@@
Care is required in specifying what is meant here by ‘horizon’, since there are several such notions (event horizon, trapping horizon, apparent horizon, dynamical horizon…) which in this context may give tiny differences in location. For precision, by ‘horizon’ we mean here the event horizon, if this exists. If it doesn’t, we mean the boundary of the past of a late-time spacelike region lying outside the black hole (say outside the trapping region). With this definition, the horizon is light-like.
@@
Now consider the evolution of the quantum fields from $\Sigma_{1}$ to $\Sigma_{2}$. We are in a region far away from the singularity and therefore (assuming the black hole is large) from high curvature. Therefore we expect conventional quantum field theory to hold here, without strange quantum gravity effects, at least up to high energy scales. Since the horizon is light-like, $\Sigma_{1}^{in}$ is in the causal past of $\Sigma_{2}^{in}$. This implies that any local observable on $\Sigma_{1}^{in}$ is fully determined by observables on $\Sigma_{2}^{in}$. That is, if ${\cal A}_{i}$ is the local algebra of observables on $\Sigma_{i}^{in}$ then ${\cal A}_{1}$ is a subalgebra of ${\cal A}_{2}$:
@@
Therefore any state on ${\cal A}_{2}$ is also a state on ${\cal A}_{1}$ and if two such states can be distinguished by observables in ${\cal A}_{1}$ they certainly can be distinguished by observables in ${\cal A}_{2}$ as the former are included in the latter. Therefore the states that can be distinguished by ${\cal A}_{1}$ —which is to say: on $\Sigma_{1}^{in}$— can also be distinguished by ${\cal A}_{2}$ —which is to say: on $\Sigma_{2}^{in}$. Therefore the distinguishable states on $\Sigma_{1}^{in}$ are a subset of those in $\Sigma_{2}^{in}$. How many are they? Either there is an infinite number of them, or a finite number due to some high-energy (say Planckian) cut-off. If there is an infinite number of them, then immediately the number of states distinguishable from inside the black hole is larger that $N_{NB}$, which is finite. If there is a finite number of them, then the number $N_{2}$ of distinguishable states on $\Sigma_{2}^{in}$ must be equal or larger than the number $N_{2}$ of states distinguishable on $\Sigma_{1}^{in}$, because the second is a subset of the first. That is
@@
The physical interpretation of the conclusion is simple: the thermal behavior of the black hole described by the Bekenstein-Hawking entropy $S=A/4$ is determined by the physics of the vicinity of the horizon.
A vivid manifestation of the fact that in classical GR the effect of a black hole on its surroundings is independent of the black hole interior is in the numerical simulations of black hole merging and radiation emission by oscillating black holes: in writing the numerical code, it is routine to cut away a region inside the (trapping) horizon: it is irrelevant for whatever happens outside! This is true in classical GR, and there is no compelling reason to suppose it to fail if quantum fields are around.
@@
In conclusion, if there are more states available in a black hole than $e^{A/4}$, then the Page argument for the information loss paradox fails. States purifying the Hawking radiation are inside the hole even when the horizon shrinks. The information can later escape from the remnant emission (see below).